Circular No. 6726

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970828 R. Remillard, A. Wood, D. Smith, and A. Levine, Massachusetts In-

stitute of Technology, communicate: “The RXTE All Sky Monitor has detected the γ-ray burst of Aug. 28.73931 UT that was logged as BATSE trigger no. 6350. The burst was detected by ASM camera 1 with a mean flux of 756 mCrab (2–12 keV) averaged over the 40 s in which the burst was seen before the camera was rotated. Ninety seconds later, ASM camera 2 scanned over the source and determined a crossing line of position and an average flux of 238 mCrab, declining in an exponential-like fashion with a 30-s time constant. The x-ray emission began roughly simultaneously with the γ-ray trigger, but was detectable with the ASM over at least 160 s. The combined camera detections yield an x-ray position of α = 18^{h}08^{m}29^{s}, δ = +59^{o}18′.0 (equinox 2000.0; 1σ uncertainties of ± 3′.2 in α and ± 1′.2 in δ). Optical observations are urgently needed.”

XTE J1755–324 R. N. Ogley and T. D. C. Ash, Open Univeristy; and R. P. Fender,

Sussex University, report that a search was made on Aug. 18.17–18.69 UT for a radio counterpart to the recently detected x-ray source XTE J1755– 324 (IAUC 6710). Observations were made using the Australia Telescope Compact Array at 1384 and 2496 MHz, and no radio emission was detected down to a level of 0.3 mJy (1384 MHz) and 0.2 mJy (2496 MHz) within the 1′ error box of XTE. The nearest source detected was 7′ from the reported position of the x-ray source. At the time of these observations, XTE J1755– 324 had an x-ray (2–12 keV) flux of ∼ 7 counts/s, measured by the ASM on XTE.

COMET C/1997 P2 (SPACEWATCH) Extension to ephemeris on IAUC 6719 (elements MPEC 1997-Q02):

1997 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Aug. 20 21^{h}36.^{m}76 − 6^{o}30.^{′}6 3.197 4.202 172.^{o}2 1.^{o}9 18.8
30 21 32.67 − 6 24.0 3.216 4.198 164.6  3.7 18.8
Sept. 9 21 29.19 − 6 19.0 3.261 4.195 154.9  5.8 18.8
19 21 26.68 − 6 13.9 3.332 4.195 145.0  7.9 18.8
29 21 25.39 − 6 07.1 3.425 4.196 135.2  9.7 18.9
Oct. 9 21 25.49 − 5 57.5 3.538 4.198 125.6 11.2 19.0
19 21 27.05 − 5 43.9 3.666 4.203 116.3 12.3 19.1
1997 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green